Circular No. 6706

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997cy IN ANONYMOUS GALAXY The Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC

6639) reports the discovery of an apparent supernova (V = 17.4, R = 16.8) on CCD images taken on July 16 by S. Sabine with the Mt. Stromlo 1.27-m telescope (+ Macho Camera). SN 1997cy is located on a faint host galaxy near Abell 3266 at α = 4^{h}32^{m}54^{s}.86, δ = −61^{o}42′57′′.5 (equinox 2000.0). A nearby star has position end figures 59^{s}.27, 43′06′′.6. SN 1997cy is not visible on images taken on Mar. 12 (limiting mag R = 20.5). A V CCD frame taken on July 22 by D. Baines and J. Howard with the Mt. Stromlo 0.76-m telescope shows that SN 1997cy had faded by 0.5 mag.

SNR G327.6+14.6 T. Tanimori, Y. Hayami, and S. Kamei, Tokyo Institute of Technology,

Tokyo; and T. Kifune, ICRR, University of Tokyo, report on behalf of the CANGAROO Collaboration: “We have detected TeV γ-rays from the remnant of the supernova of 1006 (cf. Green 1984, MNRAS 209, 449) with a statistical significance of 5.5σ, in data taken over 26 hr in April and June 1996 with the 3.8-m imaging Air Cherenkov Telescope at Woomera. It is the first evidence of very-high-energy γ-rays from an extended supernova remnant. We have an additional observation (6.0σ significance) in March and April 1997 to confirm the result. The TeV γ-rays are mainly from the northeast rim of the shell of the supernova remnant, in accordance with the nonthermal x-rays (2–10 keV) that the x-ray satellite ASCA has detected. The γ-ray signal suggests extended emission along the ridge of the northeast rim. However, by approximating the emission as from a single point source at the rim, preliminary estimation of the integral flux for the 1996 data is (3.0 ± 0.54) × 10^{−12} cm^{−2} s^{−1} (> 3 TeV). The flux will be increased by taking into account the extended emission that is much wider than the point-spread function of our detection (∼ 0^{o}.2 in FWHM).”

COMET C/1995 O1 (HALE-BOPP) Total visual magnitude estimates: July 9.82 UT, 2.9 (D. A. J. Seargent,

The Entrance, N.S.W., 25×100 binoculars); 10.92, 3.3 (A. Pearce, Subiaco, W. Australia, 20×80 binoculars); 11.92, 3.6 (Pearce); 14.92, 3.8 (Pearce); 16.82, 3.1 (Seargent, naked eye); 16.92, 3.9 (Pearce).

1997 July 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green